describe the structure of the catalytic domain of transpeptidase in terms of the numbers and types of the various secondary structure elements present and the way these are arranged to give tertiary structure
Answer #1

Answer

Transpeptidase may refer to DD-transpeptidase, a bacterial enzyme that cross links the peptidoglycan chains to form rigid cell walls. A protein-sorting transpeptidase (e.g. sortase), that cleaves a C- terminal sorting signal from its target proteins and then covalently attaches the remainder to the cell surface.

Crosslinking of peptidyl moieties of adjacent glycan is a two-step reaction. The first step involves the cleavage of the D-alanyl-D-alanine bond of a peptide unit precursor acting as carbonyl donar, the release of the carboxyl-terminal D-alanine and the formation of the acyl-enzyme intermediate and the formation of a new peptide bond between the carbonyl of the D-alayl moiety and the amino groupe of another peptide unit.

Most discussion of DD-peptidase mechanisms revolves around the catalysts of proton transfer. During formation of the acyl-enzyme intertmediate, a proton must be removed from the active site serine hydroxyl group and one must be added to the amine leaving group. A similar proton movement must be facilitated in deacylation. The identity of the general acid and base catalysts involved in these proton transfers has not yet been elucidated. However, the catalytic triad tyrosine, lysine and serine as well as serine, lysine serine have been proposed.

Transpeptidase are members of the penicilloyl-serine transferase superfamily, which has a signature SxxK conserved motif. With x denoting a variable amino acid residue, the transpeptidases of this superfamily show a trend in the form of three motifs: Sxxk, SxN and KTG. These motifs occur at equivalent places and are roughly equally spaced, along the polypeptide chain. The folded protein brings these motifs close to each other at the catalytic center between an all \alpha domain and an \alpha /\beta domain.

The structure of the streptomyces K15 DD-transpeptidase has been studied, and consist of a single polypeptide chain organized into two domains. One domain contains mainly \alpha -helices, and the second one is of \alpha /\beta -type. The center of the catalytic cleft is occupied by the Ser35-Thr36-Thr37-Lys38 tetrad, which includes the nuclephilic Ser35 residue at the amino-terminal end of helix \alpha 2. One side of the cavity is defind by the Ser96-Gly97-Cys98 loop connecting helices \alpha 4 and \alpha 5. The Lys213-Thr214-Gly215 triad lies on strand \beta 3 on the opposite side of the cavity. The backbone NH group of the essential Ser35 residue and that of Ser216 downstream from the motif Lys213-Thr-214-Gly215 occupy positions that are compatible with the oxyanion hole function required for catalysis.

The enzyme is classified as a DD- transpeptidase because the susceptible peptide bond of the carbonyl donor extends between two carbon atoms with the D-configuration.